# TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning

### v2

@@ B.2 Term Explanation
Criteria
Description
Cell Position
Specify the range or position of the cells you want to search. For example, you may want to search for explanations only in the cells of a specific column, row, or a particular section of the table.
Cell Content
Define the specific content or data type within the cells you want to search. For instance, you may want to search for explanations in cells containing numerical values, dates, specific keywords, or a combination of certain words.
Cell Formatting
Consider the formatting or styling applied to the cells. This could include searching for explanations in cells with bold or italic text, specific background colors, or cells that are merged or highlighted in a certain way.
Cell Context
Take into account the context surrounding the cells. You can search for explanations in cells that are adjacent to certain labels, headings, or identifiers, or within a specific context provided by other cells in the same row or column.
Cell Properties
Consider any specific properties associated with the cells. This might include searching for explanations in cells that have formulas, links, or other data validation rules applied to them.
@@ Appendix D Additional Experiment Settings
Table Reasoning Tasks. Each instance in table-based reasoning consists of a table $T$, a natural language question $Q$, and an answer $A$. Specifically, table $T$ is defined as $T=\left\{v_{i,j}\mid i\leq\operatorname{R}_{T},j\leq\operatorname{C}_{T}\right\}$, containing $R_{T}$ rows and $C_{T}$ columns. The content of the cell in the $i$-th row and $j$-th column is represented by $v{i,j}$. A question $Q$ is a sequence of $n$ tokens: $Q=\{q_{1},q_{2},q_{3},\cdots,q_{n}\}$. In this paper, our primary focus is on two distinct table-based reasoning tasks, table-based fact verification (TFV) and table-based question answering (TQA). In TFV, the answer $A$ is a boolean value in $\{0,1\}$, indicating the veracity of the input statement (where $1$ means the statement is entailed by the given table, and 00 means the statement is refuted by the given table). In TQA, the answer is a sequence of natural language tokens represented as $A=\{a_{1},a_{2},a_{3},\cdots,a_{n}\}$ corresponding to the posed question. For our experiments, all tables first undergo table sampling and table augmentation by our proposed method and then are serialized into a sequence by table packing and serialization. Note that various serialization methods have already been explored in our previous work Sui et al. (2023). Detailed implementation specifics are provided in Section §2.3.
@@ D.1 Downstream Tasks and Datasets
Dimension/Measure
Metadata-based
Distinguish each element in a table as either dimension field or measure field.
Semantic Field Type
Metadata-based
Classify the meaning and format of the data within each field based on knowledge graphs.
Table Size
Metadata-based
Basic information of a table including numbers of rows and columns.
Statistics Feature
Metadata-based
Statistics features such as change rate, numerical distribution, range of data.
Header Hierarchy
Metadata-based
The organization and structure of header elements within a table.
Docs References
Retrieval-based
External domain knowledge from reliable webpages (e.g., wikipedia, Wolfram Alpha, etc.) which are similar to the given context.
Term Explanation
Retrieval-based
External domain knowledge such as term and metric definitions (formulas, relevant documents/sources, search results, etc.)
Self Prompting
Self-consistency-based
Leverage LLMs to generate some reasoning thoughts as supplementary for table augmentation (self-augmented prompting, chain-of-thoughts, etc.)
@@
Metrics. For TQA and TFV tasks (SQA, FEVEROUS, TabFact and HybridQA), we report the exact match accuracy of answer sets. For data-to-text generation task (ToTTo), we report BLEU-4 score.
@@ D.2 Models
In this study, we evaluate the performance of the recent dominant LLM models, 1) Instruct-GPT-3.5 Ouyang et al. (2022), using versions gpt-3.5-turbo, gpt-3.5-turbo-16k; 2) GPT-4, using the latest version of gpt-4 model; 3) Llama-2-70B  Touvron et al. (2023), using version 17; 4) Mixtral-8x7B  Jiang et al. (2024), using version 0.1.
Unless otherwise specified, we utilize gpt-3.5-turbo in all experiments. In the sampling methods, we use text-embedding-ada-002 ope for row and query embedding generation. The comparison experiments using other embeddings models, such as, text-search-ada-doc-001, bge-largen-en Xiao et al. (2023), all-MinLM-L6-v2 Reimers and Gurevych (2019) can be found in Table 7. The development of TAP4LLM begins with the foundation provided by LLMs. In designing our framework, we opt to use OpenAI models as our base model due to their excellent capabilities in language reasoning. However, the choice is not exclusive. Since TAP4LLM use natural language as an intermediary for interactive communication between the table and LLMs, it can also support other outstanding open-sourced models using natural language as input, such as Phoenix Chen et al. (2023), ChatGLM Zeng et al. (2022), Ziya IDEA-CCNL (2023), and Baichuan Intelligence (2023). This design provides versatility and flexibility in TAP4LLM implementation.
@@ E.3 TAP4LLM in Open-source model
Beyond conducting experiments on GPT models. we also evaluate the effectiveness of TAP4LLM on two most updated LLMs: Llama-2-70B and Mixtral-8x7B. According to Table 10, we first evaluated direct inference on open-source models and then apply TAP4LLM to each model. The result demonstrates that TAP4LLM increases models’ performance on all five datasets. The experiment gives us insights on the advantages of TAP4LLM among general LLMs. We will also conduct experiments on other table-related pre-trained models e.g.TaBERT Yin et al. (2020) in our future work.
@@ E.4 TAP4LLM in Database Application
Dataset. We test TAP4LLM effectiveness on Spider Yu et al. (2018). Spider is a cross-domain Text-to-SQL dataset as shown in Table 5. Each instance contains a natural language question, a specific database containing tabular information, and one corresponding SQL query.
Metric. We evaluate TAP4LLM on the development split Spider-dev which contains 1034 instances over 200 databases. We use the Execution Accuracy, followed by the original paper Yu et al. (2018), to compare the execution output of the predicted SQL query with golden SQL query.
@@
Table 12: Comparative results of table augmentation methods on Spider. We use semantic-based sampling method without augmentation as the baseline for table augmentation.

### v3

@@ B.2 Term Explanation
Criteria Description
Cell Position Specify the range or position of the cells you want to search. For example, you may want to search for explanations only in the cells of a specific column, row, or a particular section of the table.
Cell Content Define the specific content or data type within the cells you want to search. For instance, you may want to search for explanations in cells containing numerical values, dates, specific keywords, or a combination of certain words.
Cell Formatting Consider the formatting or styling applied to the cells. This could include searching for explanations in cells with bold or italic text, specific background colors, or cells that are merged or highlighted in a certain way.
Cell Context Take into account the context surrounding the cells. You can search for explanations in cells that are adjacent to certain labels, headings, or identifiers, or within a specific context provided by other cells in the same row or column.
Cell Properties Consider any specific properties associated with the cells. This might include searching for explanations in cells that have formulas, links, or other data validation rules applied to them.
@@ D.1 Downstream Tasks and Datasets
Table Reasoning Tasks. Each instance in table-based reasoning consists of a table $T$, a natural language question $Q$, and an answer $A$. Specifically, table $T$ is defined as $T=\left\{v_{i,j}\mid i\leq\operatorname{R}_{T},j\leq\operatorname{C}_{T}\right\}$, containing $R_{T}$ rows and $C_{T}$ columns. The content of the cell in the $i$-th row and $j$-th column is represented by $v{i,j}$. A question $Q$ is a sequence of $n$ tokens: $Q=\{q_{1},q_{2},q_{3},\cdots,q_{n}\}$. In this paper, our primary focus is on two distinct table-based reasoning tasks, table-based fact verification (TFV) and table-based question answering (TQA). In TFV, the answer $A$ is a boolean value in $\{0,1\}$, indicating the veracity of the input statement (where $1$ means the statement is entailed by the given table, and 00 means the statement is refuted by the given table). In TQA, the answer is a sequence of natural language tokens represented as $A=\{a_{1},a_{2},a_{3},\cdots,a_{n}\}$ corresponding to the posed question. For our experiments, all tables first undergo table sampling and table augmentation by our proposed method and then are serialized into a sequence by table packing and serialization. Detailed implementation specifics are provided in Section §2.3.
@@
Dimension/Measure Metadata-based Distinguish each element in a table as either dimension field or measure field.
Semantic Field Type Metadata-based Classify the meaning and format of the data within each field based on knowledge graphs.
Table Size Metadata-based Basic information of a table including numbers of rows and columns.
Statistics Feature Metadata-based Statistics features such as change rate, numerical distribution, range of data.
Header Hierarchy Metadata-based The organization and structure of header elements within a table.
Docs References Retrieval-based External domain knowledge from reliable webpages (e.g., wikipedia, Wolfram Alpha, etc.) which are similar to the given context.
Term Explanation Retrieval-based External domain knowledge such as term and metric definitions (formulas, relevant documents/sources, search results, etc.)
Self Prompting Self-consistency-based Leverage LLMs to generate some reasoning thoughts as supplementary for table augmentation (self-augmented prompting, chain-of-thoughts, etc.)
@@
Metrics. For TQA and TFV tasks (SQA, FEVEROUS, TabFact and HybridQA), we report the exact match accuracy of answer sets. For the data-to-text generation task (ToTTo), we report the BLEU-4 score.
@@ D.2 Models
We evaluate the performance of the recent dominant LLM models, 1) Instruct-GPT-3.5 Ouyang et al. (2022), using versions gpt-3.5-turbo, gpt-3.5-turbo-16k; 2) GPT-4, using the latest version of gpt-4 model; 3) Llama-2-70B  Touvron et al. (2023), using version 17; 4) Mixtral-8x7B  Jiang et al. (2024), using version 0.1.
Unless otherwise specified, we utilize gpt-3.5-turbo in all experiments. In the sampling methods, we use text-embedding-ada-002 ope for row and query embedding generation. The comparison experiments using other embeddings models, such as, text-search-ada-doc-001, bge-largen-en Xiao et al. (2023), all-MinLM-L6-v2 Reimers and Gurevych (2019) can be found in Table 8. We set the temperature of all the models as 0, top p as 1.0, frequency penalty as 0, and presence penalty as 0.
The development of TAP4LLM begins with the foundation provided by LLMs. In designing our framework, we opt to use OpenAI models as our base model due to their excellent capabilities in language reasoning. However, the choice is not exclusive. Since TAP4LLM use natural language as an intermediary for interactive communication between the table and LLMs, it can also support other outstanding open-sourced models using natural language as input, such as Phoenix Chen et al. (2023), ChatGLM Zeng et al. (2022), Ziya IDEA-CCNL (2023), and Baichuan Intelligence (2023). This design provides versatility and flexibility in TAP4LLM implementation.
@@ E.3 TAP4LLM in Open-source model
Beyond conducting experiments on GPT models. we also evaluate the effectiveness of TAP4LLM on two recent LLMs: Llama-2-70B and Mixtral-8x7B. According to Table 11, we first evaluated direct inference on open-source models and then apply TAP4LLM to each model. The result demonstrates that TAP4LLM increases models’ performance on all five datasets.
In addition, TAP4LLM is currently evaluated under the setting of in-context learning. However, through parameter-efficient fine-tuning Hu et al. (2021) and recently advanced prompt compression techniques Jiang et al. (2023a); Zou et al. (2024), we can directly apply TAP4LLM on more challenging tabular reasoning tasks requiring training procedures. We will leave this as our future work.
@@ E.4 TAP4LLM in Database Application
Dataset We test TAP4LLM effectiveness on Spider Yu et al. (2018). Spider is a cross-domain Text-to-SQL dataset as shown in Table 6. Each instance contains a natural language question, a specific database containing tabular information, and one corresponding SQL query.
Metric We evaluate TAP4LLM on the development split Spider-dev which contains 1034 instances over 200 databases. We use the Execution Accuracy, followed by the original paper Yu et al. (2018), to compare the execution output of the predicted SQL query with the golden SQL query.
@@
Table 13: Comparative results of table augmentation methods on Spider. We use semantic-based sampling method without augmentation as the default method for table augmentation.